OpenStudy (anonymous):

Suppose a lab needs to make 400 liters of a 32% acid solution, but the only solutions available to the lab are 20% acid and 65% acid. What system of equations can be used to find the number of liters of each solution that should be mixed to make the 32% solution? Let c represent the number of liters of 20% acid solution and let d represent the number of liters of 65% acid solution.

OpenStudy (nikato):

You just have to set up the equations,right?

OpenStudy (anonymous):

yes

OpenStudy (nikato):

Ok. Lets call the 20%= solution A And the 65%= solution B And the 32%= mixture

OpenStudy (nikato):

The basis setup would be (liters of A x concentration of A) + ( liters of B x concentration of B)= (liters of mixture x concentration of mixture)

OpenStudy (nikato):

And you know there are "c" liters of A with a concentration of 20% or 0.2

OpenStudy (nikato):

"d" liters of B with a concentration of 0.65

OpenStudy (nikato):

And 400 liters of the mixture with a concentration of 0.32

OpenStudy (nikato):

Plug in all the given info to the setup I gave u. 0.2c + 0.65d = 0.32(400)

OpenStudy (nikato):

Also because the mixture is 400 liters You would need this equation c+d=400

OpenStudy (anonymous):

thanks

OpenStudy (nikato):

youre welcome!